AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 161     Entry time: Tue Nov 23 08:49:33 2021
Author: CB + ML 
Subject: NEG activation completed 

Found temperatures and pressure stabilised overnight. See attached.

12:00 Began activation of 3x UHV1400 in channel 2, top module. Started using 90 W as per procedure. No effect on temperature. Increased to 120 W. OK

14:30 Began activation of 2x UHV1400 in channel 1, in front of the detectors. Considered moving detectors but decided against it since bellows are rated to move at 20C at most. Activated at 60 W (vs. 75 W in procedure). OK

 

All NEG elements off. CARME heaters ramping down to 60 C overnight. 

Ion pumps activated < 100 C. Valve closed. #4 doesn't come online properly. Sone issues with #2 as well but it recovered on its own. Left cooling down. No changes. Ion pumps off and valve open. Will try tomorrow. Could be faulty cable?

 

Vacuum division may want to activate their pump as well - times unclear.

Attachment 1: vacplot24.png  103 kB  | Hide | Hide all
vacplot24.png
Attachment 2: CARMEtemps.png  124 kB  | Hide | Hide all
CARMEtemps.png
Attachment 3: vacplot24-1.png  102 kB  Uploaded Tue Nov 23 15:44:42 2021  | Hide | Hide all
vacplot24-1.png
Attachment 4: CARMEtemps-1.png  139 kB  Uploaded Tue Nov 23 15:44:42 2021  | Hide | Hide all
CARMEtemps-1.png
ELOG V3.1.4-unknown